Please help me solve this problem
prove that
for all values ok k ,roots of this equation ( x^2-2((k)+(1/k))x+3=0) are real

Answer by oscargut(2103)

x^2-2((k)+(1/k))x+3=0
discriminant = d = (-2((k)+(1/k)))^2-4(1)(3) =
4(k^2+2+1/k^^2) -12
= 4k^2+8+4/k^2 -12
= 4k^2+4/k^2 -4
= 4(k^2+1/k^2 -1)
= 4 ((k-1/k)^2 +1) >0
Then the equation has 2 real solutions


I can help you with more questions at: mthman@gmail.com
Thanks